OpenStudy (anonymous):

Simplify each expression. −− √125

OpenStudy (anonymous):

5sqrt5

OpenStudy (anonymous):

\[\sqrt{125} = \sqrt{5^3} = 5^{3/2} = 5^{2/2}*5^{1/2} = 5*\sqrt{5}\]